OpenStudy (anonymous):

derivative of (t-1/t)

OpenStudy (anonymous):

For t!=0, f(t) = 1-(1/t). Do you know how to derive that?

OpenStudy (anonymous):

Do you mean \[\frac{t-1}{t}\] or \[t-\frac{1}{t}\]?

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

Oh nvm forget my answer then.

OpenStudy (anonymous):

Well you're just doing two separate derivatives then. One for t, and one for -1/t. Can you do these separately?

OpenStudy (anonymous):

no i wouldn't have asked othewrwise

OpenStudy (anonymous):

i think the derivative of t is zero

OpenStudy (anonymous):

You don't know what the derivative of t is?

OpenStudy (anonymous):

or 1

OpenStudy (anonymous):

No, 0 is only the derivative of a constant. For example, the derivative of the number 5 is 0. t is a variable that changes, so its derivative can't be 0 (that would imply it's not changing). Yes, 1 is correct for t.

OpenStudy (anonymous):

now -1/t?

OpenStudy (anonymous):

Use to power rule: \[\frac{d}{dx}x^n = n\cdot x^{n-1}\] In the first case, n = 1, and in the second case n = -1.

OpenStudy (anonymous):

If you know the power rule, you can use it to derive -1/t by first rewriting it as \[-t^{-1}\]

OpenStudy (anonymous):

thank wio